# Fully Open Source Moxin-LLM Technical Report

### v2

@@ Abstract
Recently, Large Language Models (LLMs) have undergone a significant transformation, marked by a rapid rise in both their popularity and capabilities. Leading this evolution are proprietary LLMs like GPT-4 and GPT-o1, which have captured widespread attention in the AI community due to their remarkable performance and versatility. Simultaneously, open-source LLMs, such as LLaMA and Mistral, have made great contributions to the ever-increasing popularity of LLMs due to the ease to customize and deploy the models across diverse applications. Although open-source LLMs present unprecedented opportunities for innovation and research, the commercialization of LLMs has raised concerns about transparency, reproducibility, and safety. Many open-source LLMs fail to meet fundamental transparency requirements by withholding essential components like training code and data, and some use restrictive licenses whilst claiming to be “open-source,” which may hinder further innovations on LLMs. To mitigate this issue, we introduce Moxin 7B, a fully open-source LLM developed in accordance with the Model Openness Framework (MOF), a ranked classification system that evaluates AI models based on model completeness and openness, adhering to principles of open science, open source, open data, and open access. Our model achieves the highest MOF classification level of “open science” through the comprehensive release of pre-training code and configurations, training and fine-tuning datasets, and intermediate and final checkpoints. Experiments show that our model achieves superior performance in zero-shot evaluation compared with popular 7B models and performs competitively in few-shot evaluation.
Homepage: https://github.com/moxin-org/Moxin-LLM
@@
Chat model: https://huggingface.co/moxin-org/moxin-chat-7b
@@ 1 Introduction
The field of natural language processing has witnessed the most exciting discoveries of the last ten years with the emergence of large language models (LLMs). At the forefront of this evolution are LLMs such as GPT-4 [1], Claude [2], and Gemini [3], which have captured the attention of the AI community due to their performance and versatility. Meanwhile, the recent emergence of openly accessible yet highly capable LLMs such as LLaMA [4], Falcon [5], and Mistral [6] allow researchers and practitioners to easily obtain, customize, and deploy LLMs in more various environments and for more diverse use cases. The trends have made people eagerly asking about what’s next and some suggest “a general intelligence” is right around the corner.
@@
Despite the growing influence and accessibility of open-source LLMs, a notable challenge emerged: many model producers restrict visibility and access to their training, fine-tuning, and evaluation processes, including crucial components such as their training code and data [7]. Some model producers even use restrictive licenses whilst claiming to be “open-source.” This practice creates barriers for the broader AI research community to study, replicate, and innovate upon advanced LLMs. In parallel, it prevents businesses from fully leveraging open-source models for innovative industrial applications, as its commercialization has raised concerns about transparency, reproducibility, and safety.
@@
To unlock the full potential of LLMs and open innovation, we must return to democratize this research by putting the model into the hands of more researchers and making the datasets the models train on fully open-source. This requires moving beyond the simple sharing of model weights to embrace complete transparency in training, datasets, and implementation detail, which is crucial for fostering a more inclusive and collaborative research environment that can sustain a healthy open-source ecosystem [8].
@@
To achieve this goal, we introduce Moxin 7B, a fully open-source LLM developed by complying with the Model Openness Framework (MOF) introduced by [9]. The MOF provides a systematic ranking classification system to rate AI models based on their completeness and openness, incorporating the principles of open science, open source, open data, and open access. By promoting transparency and reproducibility, the MOF serves as a crucial tool to combat “openwashing” practices and establishes completeness and openness as primary criteria alongside the core tenets of responsible AI. Wide adoption of the MOF will cultivate a more open AI ecosystem, benefiting research, innovation, and adoption of state-of-the-art models.
@@
Our open-source LLM has released pre-training code and configurations, training and fine-tuning data, and intermediate and final checkpoints, aiming to make continuous commitments to fully open-source LLMs. Our model achieves the highest MOF classification level of “open science.” It is noteworthy that this commitment to openness has not compromised performance: our base model achieves superior performance in zero-shot evaluation compared with popular 7B models and performs competitively in few-shot evaluation. Remarkably, our chat model can outperform 7B baselines like Llama2-7B-chat. Our homepage is https://github.com/moxin-org/Moxin-LLM.
@@ 2.1 Models, Tokenizers, and Training
Models. State-of-the-art large language models (LLMs) typically comprise a substantial number of parameters, often approaching or exceeding 100 billion [4, 1, 3]. To facilitate broader accessibility, smaller models with fewer than 20 billion parameters, and even those around 7 billion parameters, have been developed [10, 11, 4, 6, 12, 13]. In addition, efficiency-enhancing techniques, such as implementing MAMBA-based architectures in Jamba, have been employed to optimize performance [12, 13].
@@
Training. To enhance the performance of smaller models beyond their inherent limitations, various training strategies can be employed. A notable example is the application of Mixture of Experts (MoE) training, which has achieved significant success in models like Mixtral [23].
@@ 2.3 Open-source datasets
3 Model Training
@@ 3.1 Model Architecture
We opt to extend the Mistral model architecture [6] due to its ability to achieve high performance while maintaining efficient inference speeds. The original Mistral 7B model demonstrates superior performance compared to multiple 7B language models and even outperforms larger models on various evaluation benchmarks. Notably, it surpasses the LLaMA 34B model [47] in tasks such as mathematics and code generation.
@@
The original Mistral model leverages grouped-query attention (GQA)[48] and sliding window attention (SWA)[49]. GQA reduces memory requirements during decoding, allowing for larger batch sizes and higher throughput, and it significantly accelerates inference speed—an essential factor in real-time applications. Meanwhile, SWA effectively handles long sequences without incurring substantial computational overhead. By incorporating these techniques, the model achieves significant improvements in performance and efficiency, which we have adopted in our extended model.
Building upon the original Mistral model, which consists of 32 blocks, we have extended the architecture to 36 blocks. Furthermore, we also employ GQA to partition the query heads into multiple groups, each sharing a single key head and value head. This approach interpolates between multi-query attention (MQA) and multi-head attention (MHA) in large language models, striking a balance between the computational speed of MQA and the representational quality of MHA, thereby providing a favorable trade-off. Additionally, our model incorporates a rolling buffer cache with a fixed attention span, effectively limiting cache size and preventing excessive memory usage when processing long sequences.
3.2 Training Data
3.2.1 Text Data
@@ 3.3 Training Configuration
The total number of tokens used for pre-training our Moxin-7B model is over 2T, and the pre-training process consists of three phases. In the first phase, we use pre-training corpora with the context length of 2k. In the second phase, we use pre-training corpora with the context length of 4k. In the third phase, we utilize the capability-specific enhancement data. We provide the model performance with only the first two phases and also with all three phases to validate the performance of the third phase.
We use Colossal-AI [68] as our training framework. Colossal-AI is a unified deep learning system that provides the fullest set of acceleration techniques for the AI community. With its modular design, ColossalAI allows for a free combination of these techniques to achieve the best training speedup. Colossal-AI’s optimized parallelism and heterogeneous training methods are employed to achieve superior system performance compared to baseline systems. These methods are provided through user-friendly APIs, requiring minimal code modifications.
@@
During training, AdamW [69] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-8}$ and weight decay = 0.1 is used to optimize the model. We use the cosine learning rate decay and the learning rate decays to 10% of its maximum. Learning Rate is set to $2e^{-6}$.
3.4 Alignment
Following the pre-training phase, we fine-tune the model into a helpful and harmless AI assistant. In our Alignment stage, we mainly use supervised fine-tuning (SFT), during which we fine-tune the model to follow diverse human instructions by high-quality instruction data. We use the Tulu v2 dataset [70] for instruction tuning. The dataset consists of a mix of FLAN, Open Assistant 1, ShareGPT, GPT4-Alpaca, LIMA, and so on.
3.5 Long-Context
To deal with the long-context problem, our model leverages grouped-query attention (GQA) [48], sliding window attention (SWA) [49], and Rolling Buffer Cache [6]. GQA reduces the memory requirement during decoding, allowing for higher batch sizes hence higher throughput.
Besides, SWA can handle longer sequences more effectively at a reduced computational cost, thereby alleviating a common limitation in LLMs. SWA exploits the stacked layers of a transformer to attend information beyond the window size $W$. At the last layer, with SWA, using a window size of $W=4096$, we have a theoretical attention span of approximately $14K$ tokens or above.
Our model adopts Rolling Buffer Cache which limits the cache size using a rolling buffer cache with a fixed attention span. The cache has a fixed size of $W$, and the keys and values for the timestep $i$ are stored in position $i$ mod $W$ of the cache. As a result, when the position $i$ is larger than $W$, past values in the cache are overwritten, and the size of the cache stops increasing. On a sequence length of $32k$ tokens, this reduces the cache memory usage by 8$\times$, without impacting the model quality.

### v3

# 7B Fully Open Source Moxin-LLM – From Pretraining to GRPO-based Reinforcement Learning Enhancement

@@ Abstract
Recently, Large Language Models (LLMs) have undergone a significant transformation, marked by a rapid rise in both their popularity and capabilities. Leading this evolution are proprietary LLMs like GPT-4 and GPT-o1, which have captured widespread attention in the AI community due to their remarkable performance and versatility. Simultaneously, open-source LLMs, such as LLaMA and Mistral, have made great contributions to the ever-increasing popularity of LLMs due to the ease to customize and deploy the models across diverse applications. Although open-source LLMs present unprecedented opportunities for innovation and research, the commercialization of LLMs has raised concerns about transparency, reproducibility, and safety. Many open-source LLMs fail to meet fundamental transparency requirements by withholding essential components like training code and data, and some use restrictive licenses whilst claiming to be “open-source,” which may hinder further innovations on LLMs. To mitigate this issue, we introduce Moxin 7B, a fully open-source LLM developed in accordance with the Model Openness Framework (MOF), a ranked classification system that evaluates AI models based on model completeness and openness, adhering to principles of open science, open source, open data, and open access. Our model achieves the highest MOF classification level of “open science” through the comprehensive release of pre-training code and configurations, training and fine-tuning datasets, and intermediate and final checkpoints, aiming to make continuous commitments to fully open-source LLMs. The pre-training cost is about \$160,000. After pre-training and obtaining the base model, we finetune the Moxin Base model with SOTA post-training framework and instruction data to obtain Moxin Instruct model. To improve the reasoning capability, we further finetune our Instruct model with chain-of-thought data distilled from DeepSeek R1, and then use Group Relative Policy Optimization (GRPO), an efficient and effective reinforcement learning algorithm following DeepSeek R1, to finetune our model, leading to the Moxin Reasoning model. Experiments show that our models achieve superior performance in various evaluations such as zero-shot evaluation, few-shot evaluation, and CoT evaluation, compared with other base models and instruct models. The outstanding performance of our Reasoning model demonstrates the effectiveness of RL for small LLMs such as 7B models. Besides our open pretraining with released base model, data, code, etc., in our post-training including the instruction finetuning and CoT finetuning, we adopt open-source training frameworks with available data, code, and configurations. We release our instruct model and reasoning model, along with the available data and code to derive these models.
Homepage with all codes: https://github.com/moxin-org/Moxin-LLM
@@
Instruct model: https://huggingface.co/moxin-org/moxin-Instruct-7b
Reasoning model: https://huggingface.co/moxin-org/moxin-Reasoning-7b
@@ 1 Introduction
The field of natural language processing has witnessed the most exciting discoveries of the last ten years with the emergence of large language models (LLMs). At the forefront of this evolution are LLMs such as GPT-4 [1], Claude [2], and Gemini [3], which have captured the attention of the AI community due to their performance and versatility. Meanwhile, the recent emergence of openly accessible yet highly capable LLMs such as LLaMA [4], Falcon [5], and Mistral [6] allow researchers and practitioners to easily obtain, customize, and deploy LLMs in more various environments and for more diverse use cases. The trends have people eagerly asking about what’s next and some suggest “a general intelligence” is right around the corner.
@@
Despite the growing influence and accessibility of open-source LLMs, a notable challenge has emerged: many model producers restrict visibility and access to their training, fine-tuning, and evaluation processes, including crucial components such as their training code and data [7]. Some model producers even use restrictive licenses whilst claiming to be “open-source.” This practice creates barriers for the broader AI research community to study, replicate, and innovate upon advanced LLMs. In parallel, it prevents businesses from fully leveraging open-source models for innovative industrial applications, as its commercialization has raised concerns about transparency, reproducibility, and safety.
@@
To unlock the full potential of LLMs and open innovation, we must democratize this research by putting models into the hands of more researchers and making the datasets the models train on fully open-source. This requires moving beyond the simple sharing of model weights to embrace complete transparency in training, datasets, and implementation detail, which is crucial for fostering a more inclusive and collaborative research environment that can sustain a healthy open-source ecosystem [8].
@@
To achieve this goal, we introduce Moxin 7B, a fully open-source LLM developed by complying with the Model Openness Framework (MOF) introduced by [9]. The MOF provides a systematic ranking classification system to rate AI models based on their completeness and openness, incorporating the principles of open science, open source, open data, and open access. By promoting transparency and reproducibility, the MOF serves as a crucial tool to combat “openwashing” and to establish completeness and openness as primary criteria alongside the core tenets of responsible AI. The wide adoption of MOF and open-source state-of-the-art models will cultivate a more open AI ecosystem, benefiting research and innovation.
@@
Our open-source LLM has released pre-training code and configurations, training and fine-tuning data, and intermediate and final checkpoints, aiming to make continuous commitments to fully open-source LLMs. Our model achieves the highest MOF classification level of “open science”. It is noteworthy that this commitment to openness has not compromised performance: our base model achieves superior performance in zero-shot evaluation compared to popular 7B models and performs competitively in few-shot evaluation.
We also finetune the Moxin Base model with SOTA post-training framework and instruction data to obtain Moxin Instruct model. To improve the reasoning capability, we further finetune our model with chain-of-thought data distilled from DeepSeek R1, and then use Group Relative Policy Optimization, an efficient and effective reinforcement learning algorithm following DeepSeek R1, to finetune our model, leading to the Moxin Reasoning model. Experiments show that our models achieve superior performance in various evaluations such as zero-shot evaluation, few-shot evaluation, and CoT evaluation, compared with other instruct and reasoning models. The outstanding performance of our Reasoning model demonstrates the effectiveness of RL for small LLMs such as 7B models.
Besides our open pretraining with released base model, data, code, etc., in our post-training including the instruction finetuning and CoT finetuning, we adopt open-source training frameworks with available data, code, and configurations. We release our instruct model and reasoning model, along with the available data and code to derive these models. Our homepage is https://github.com/moxin-org/Moxin-LLM. We summarize our contributions below.
  • For the pre-training, we release the Moxin Base 7B model, together with all training code, data, and checkpoints, aiming to make continuous commitments to fully open-source LLMs. Our base model performs competitively compared with SOTA pre-trained base models, with a moderate training cost about \$160,000.
  • To further improve the model performance, we post-train our base model with SFT and DPO methods, based on the open-source Tülu 3 framework. Two versions of instruction models are developed with two different open-source datasets (Tülu 3 and Infinity Instruct), respectively. Both models are available and the one finetuned with the Tülu 3 dataset is released as Moxin Instruct 7B, with superior performance in different evaluations.
  • To improve the reasoning capabilities of the model with CoT, we further finetune our instruct model with SFT based on reasoning data from Open-Thoughts and OpenR1-Math-220k, and then reinforcement learning (GRPO following DeepSeek R1). We develop two versions of reasoning models with two different open-source RL frameworks (DeepScaleR and AReaL) and their corresponding data, respectively. Both models are available and the one finetuned with DeepScaleR is released as Moxin Reasoning 7B, with superior performance in CoT evaluations. We demonstrate that RL such as GRPO can work effectively for small LLMs.
We summarize our open-source releases below.
  • Pre-training code, data, and Moxin Base model.
  • Post-training code, data, and Moxin Instruct model.
  • RL code, data and Moxin Reasoning model.
@@ 2.1 Models, Tokenizers, and Training
Models. State-of-the-art large language models (LLMs) typically comprise a substantial number of parameters, often approaching or exceeding 100 billion [4, 1, 3]. To facilitate broader accessibility, smaller models with fewer than 20 billion parameters, and even those around 7 billion parameters, have been developed [10, 11, 4, 6]. In addition, efficiency-enhancing techniques, such as pruning [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], token reduction [25, 26] or implementing MAMBA-based architectures in Jamba, have been employed to optimize performance [27, 28].
@@
Training. To enhance the performance of smaller models beyond their inherent limitations, various training strategies can be employed. A notable example is the application of Mixture of Experts (MoE) training, which has achieved significant success in models such as Mixtral [38].
@@ 2.3 Open-source datasets
Besides model pre-training, post-training has become a vital step in refining behaviors and unlocking new capabilities in LLMs. Post-training includes multiple techniques such as instruction tuning and reinforcement learning from human feedback. The sophistication and complexity of post-training approaches have continued to increase, with multiple rounds of training, model merging, leveraging synthetic data, and various training algorithms and objectives. Multiple post training datasets are released, such as Tülu 3 [62] and Infinity Instruct [63]. Tülu 3 [62] creates new datasets & new training procedures, and introduces new methods for training directly on verifiable problems with reinforcement learning. It includes both SFT training datasets and reinforcement learning datasets, together with the SFT, DPO, and PPO training recipes. Infinity Instruct [63] aims to develop a large-scale, high-quality instruction dataset with millions of high-quality instruction samples.
Furthermore, as the breakthrough of DeepSeek R1 [64], a line of SFT datasets and reinforcement learning datasets have been developed to enhance the reasoning capabilities of LLMs. Openthoughts [65] and OpenR1-Math-220k [66] datasets are obtained by feeding questions to DeepSeek R1 and collecting the reasoning responses from R1. Openthoughts [65] contains open synthetic reasoning dataset with 114k high-quality examples covering math, science, code, and puzzles. OpenR1-Math-220k [66] is a large-scale dataset for mathematical reasoning, with 220k math problems and their corresponding solutions generated by DeepSeek R1. Besides, multiple datasets are developed and released for the reinforcement learning of LLMs following DeepSeek R1, such as DeepScaleR [67] and AReaL [68, 69].
2.4 Open-source LLMs
Large language models (LLMs) have rapidly evolved into a diverse ecosystem that spans closed-source, open-weight, and fully open-sourced paradigms. On one end of the spectrum, closed-source models such as GPT-4 [1], and Gemini [70] have set high performance standards but remain accessible primarily through API services, limiting insight into their underlying architectures and training methodologies. In contrast, open-weight LLMs—while sharing their final model architectures and weights—often leave training data and many training details undisclosed. This category includes influential models Llama [71], Mistral [6], Gemma [56], Qwen [10], DeepSeek [64, 72], Baichuan [73], Phi [74], etc. Pushing the frontier further, fully open-sourced LLMs have begun to provide not only complete model weights and architectures but also the training code and datasets necessary for reproducibility. Exemplars of this fully open paradigm include Pythia [75], GPT-NeoX [76], OpenLLaMA, StarCoder [61], OLMo [77], Amber and Crystal from LLM360 [78], etc. Collectively, this vibrant landscape highlights an ongoing shift toward more accessible and fully reproducible LLMs, fostering collaboration and innovation across both academia and industry.
3 Model Pre-Training
In this section, we provide an in-depth discussion of the model training process, covering the architectural design, data curation and preprocessing, training configurations, fine-tuning for alignment, and techniques for handling long-context sequences. We detail both our methodological choices and the technical challenges addressed during the development of Moxin 7B.
@@ 3.1 Model Architecture
We extend the Mistral model architecture [6] to balance high performance with efficient inference. The original Mistral 7B model demonstrates superior performance compared to multiple 7B language models and even outperforms larger models on various evaluation benchmarks. Notably, it surpasses the LLaMA 34B model [79] in tasks such as mathematics and code generation. Key architectural enhancements include:
  • Depth Extension: While the original Mistral model comprises 32 blocks, our extended model uses 36 transformer blocks. This additional depth was empirically validated to improve both learning capacity and performance on complex downstream tasks.
  • Layer Normalization and Initialization: Pre-layer normalization is applied to stabilize the training process, accompanied by a custom initialization scheme to mitigate gradient vanishing/exploding issues in deep transformer networks.
  • Regularization Techniques: Dropout with a rate of 0.1 is employed across attention and feed-forward layers, and label smoothing is used during fine-tuning to further regularize the model.
  • Mixed-Precision and Activation Checkpointing: Mixed-precision training (FP16) is leveraged to double throughput and reduce memory consumption. Activation checkpointing is used to trade additional computation for a lower memory footprint, enabling the training of deeper networks on the available hardware.
The choice of hyperparameters, as listed in Table 1, was informed by preliminary ablation studies and a grid search over the parameter space.
3.2 Long-Context
To process extensive sequences efficiently, we integrate our long-context handling framework into the training and inference pipelines:
  • Grouped-Query Attention (GQA) [80]: We employ GQA to partition query heads into groups, where each group shares a single key and value head. This approach serves as an interpolation between multi-query attention (MQA) and multi-head attention (MHA), balancing computational speed with representational richness. GQA reduces memory requirements during decoding, enabling larger batch sizes and improved throughput.
  • Sliding Window Attention (SWA) [81]: SWA enables efficient handling of long sequences by partitioning attention into a fixed-size sliding window. This not only limits computational overhead but also preserves the ability to model dependencies beyond the immediate context window. Specifically, using a window size $W=4096$ allows the final layer to achieve a theoretical attention span exceeding 14,000 tokens.
  • Rolling Buffer Cache [6]: To further mitigate memory usage during inference, a rolling buffer cache with a fixed attention span is employed. For timestep $i$, keys and values are stored at position $i\mod W$, ensuring that once the sequence length exceeds $W$, older context tokens are overwritten rather than expanding the cache indefinitely. On sequences of 32K tokens, this approach reduces cache memory usage by approximately 8$\times$ without degrading model performance.
3.3 Training Data
3.3.1 Text Data
SlimPajama Curation and Deduplication:
@@ 3.4 Training Configuration
The pre-training of Moxin 7B spans over 2 trillion tokens and is executed in three distinct phases:
  1. 1.
    Base Pre-training: In this initial phase, the model is trained on data with a fixed context length of 2,000 tokens. This stage establishes foundational language modeling abilities, capturing syntax and basic semantic patterns.
  2. 2.
    Extended Context Pre-training: In the second phase, the context length is increased to 4,000 tokens, allowing the model to capture longer dependencies and richer context.
  3. 3.
    Capability Enhancement: The final phase incorporates the capability-specific enhancement data described earlier. We provide ablation results that compare the performance with only Phases 1 and 2 versus all three phases, thereby quantifying the benefits of the capability-enhancement data.
We use Colossal-AI [98] as our training framework with acceleration techniques including optimized parallelism, memory optimization, and heterogeneous training methods. : 1) Optimized Parallelism: We combine model, data, and pipeline parallelism to scale training across GPUs. This hybrid parallelism strategy minimizes communication overhead and ensures efficient resource utilization. 2) Memory Optimization: Techniques such as ZeRO redundancy elimination, gradient checkpointing, and mixed-precision training (FP16) are employed to maximize the effective batch size and reduce memory consumption. 3) Heterogeneous Training Methods: the dynamic scheduling and asynchronous communication protocols are leveraged to further boost training throughput. These methods are provided through user-friendly APIs, requiring minimal code modifications. The training cost is around \$160,000.
@@
During training, AdamW [99] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-8}$ and weight decay = 0.1 is used to optimize the model. We use the cosine learning rate decay, and the learning rate decays to 10% of its maximum. Learning Rate is set to $2e^{-6}$. Our micro-batch size and gradient accumulation steps were tuned to balance between memory constraints and convergence speed.
4 Model Post-Training
Following the pre-training phase, we fine-tune the model into a helpful and harmless AI assistant. We adopt the open-source Tülu 3 dataset and framework [62] for our model post-training. Tülu 3 is a family of open state-of-the-art post-trained models, alongside all of the data, data mixes, recipes, code, infrastructure, and evaluation framework, which pushes the boundaries of research in post-training. To close the performance gap between open and closed fine-tuning recipes, Tülu 3 creates new datasets & new training procedures, and introduces new methods for training directly on verifiable problems with reinforcement learning. The success of Tülu 3 is rooted in careful data curation, rigorous experimentation, innovative methodologies, and improved training infrastructure, integrating partial details from proprietary methods with novel techniques and established academic research.
For our post-training, with our base model, we follow Tülu 3 to perform supervised finetuning (SFT) and then Direct Preference Optimization (DPO). Specifically, we use the Tülu 3 SFT Mixture dataset from Tülu 3 to train our base model with the SFT training method for two epochs and obtain our SFT model, following the default training configuration of the Tülu 3 8B SFT model [62]. To promote diversity, the Tülu 3 SFT Mixture dataset contains 939,344 samples from multiple datasets such as CoCoNot [100], FLAN v2 [101], WildChat GPT-4 [102], etc., together with some data samples collected and released by the Tülu 3 team. It especially consider enhancing several capabilities that can power common use cases and the specific needs, by including following datasets: OpenMathInstruct [103] and NuminaMath [104] for mathematical reasoning, Evol-CodeAlpaca for coding, a subset of Daring-Anteater [105] for precise instruction following, SciRIFF [106] for scientific literature understanding, and TableGPT [107] for processing table-related tasks. It tracks the provenance of dataset and their subsets to verify their licenses and remove those that have issues. With thorough experimentation of Tülu 3, the final SFT data and training hyperparameters are determined to enhance target core skills without significantly impacting the performance of others, guided by the Tülu 3 evaluation framework. We use the Tülu 3 SFT Mixture dataset to train our base model with SFT for 2 epochs with a learning rate 5e-06 and a batch size of 128.
Next, we continue to train our SFT model on the Tülu 3 8B Preference Mixture dataset from Tülu 3 with the DPO training method to obtain our DPO model, following the same training configuration of the Tülu 3 8B DPO model [62]. The data creation pipeline for the Tülu 3 8B Preference Mixture dataset consists of three stages: prompt selection, response generation from a pool of models, and preference annotation with LLM-as-a-judge to create (preferred, rejected) pairs. The preference mixes data come from different prompt sources, such as SFT data, WildChat and Persona IF. It includes prompts seen during SFT training but also new, unseen prompts. To find a better training algorithm and hyper-parameters, Tülu 3 revisits the hyperparameter and algorithm choices alongside the preference datasets. It ablates both algorithm and hyperparameter choices using an early SFT checkpoint and the UltraFeedback dataset. It explores using DPO, SimPO, and length-normalized DPO, leading to the final training configuration with all hyper-parameters. We use the Tülu 3 8B Preference Mixture dataset to finetune our SFT model with DPO for 1 epochs with a learning rate 5e-07.
We also adopt the Infinity Instruct dataset [63] to train another version of our instruct model based on our base model. Infinity Instruct aims to develop a large-scale, high-quality instruction dataset. To construct a ten-million high-quality instruction dataset, it collects a large amount of open-source data as seed and iterate the dataset using two strategies: instruction selection and instruction evolution. It includes the Foundational Dataset and the Chat Dataset. The Foundational Dataset contains millions of instruction selected from open-source dataset to improve the performance of model on challenging downstream tasks (e.g., code, math). The Chat Dataset contains about 1M instructions evolved from a small subset of high-quality seed data, to further improve the instruction-following ability of model in real conversation scenarios.
5 Chain-of-Thought Enhancement
In the context of reasoning capabilities, OpenAI’s o1 series models firstly introduce inference-time scaling by increasing the length of Chain-of-Thought (CoT) reasoning process. This approach has achieved significant improvements in various reasoning tasks, such as mathematics, coding, and scientific reasoning. However, the challenge of effective test-time scaling remains an open question for the research community. DeepSeek R1 [64] takes the first step towards improving language model reasoning capabilities using pure reinforcement learning (RL). The goal is to explore the potential of LLMs to develop reasoning capabilities without any supervised data, focusing on their self-evolution through a pure RL process. After thousands of RL steps, DeepSeek-R1-Zero exhibits super performance on reasoning benchmarks.
To enhance the CoT capabilities of our model, we adopt RL techniques similar to DeepSeek R1 [64]. We first use high quality reasoning data to SFT our instruct model. The reasoning data mainly includes Openthoughts [65] and OpenR1-Math-220k [66]. These datasets are obtained by feeding questions to DeepSeek R1 and collecting the reasoning responses from R1. Openthoughts [65] contains open synthetic reasoning dataset with 114k high-quality examples covering math, science, code, and puzzles. It generates reasoning traces from DeepSeek-R1 and verifies correctness to construct the final dataset. OpenR1-Math-220k [66] is a large-scale dataset for mathematical reasoning. It consists of 220k math problems with two to four reasoning traces generated by DeepSeek R1 for problems from NuminaMath 1.5. To build OpenR1-Math-220k, it generates two answers for 400k problems using DeepSeek R1 with 512 H100 GPUs. Then it applies Math Verify to only retain problems with at least one correct answer, and leverages Llama3.3-70B-Instruct as a judge for 12% of the samples to retrieve more correct examples, leading to a filtered dataset containing 220k problems with at least one reasoning trace with a correct answer for each problem. We finetune our instruct model on these datasets with SFT based on the open-instruct finetuning framework, following the default training configuration of the Tülu 3 8B SFT model [62], with 2 epochs finetuning following a learning rate 5e-06 and a batch size of 128.
Next after SFT, we further adopt the RL techniques in DeepSeek R1, i.e., GRPO to finetune our model with RL. We adopt the DeepScaleR [67] as our RL training framework. DeepScaleR is an open-source project to fully democratize RL for LLMs and reproduce DeepSeek R1 and OpenAI O1/O3 at scale on real tasks. It opens source all training scripts (including hyperparameters), models, dataset, and logs. It is based on veRL [108], Volcano Engine Reinforcement Learning for LLM, a flexible, efficient and production-ready RL training framework designed for LLMs. veRL is flexible and easy to use with easy extension of diverse RL algorithms (combining the strengths of single-controller and multi-controller paradigms to enable flexible representation and efficient execution of complex Post-Training dataflows), seamless integration of existing LLM infra with modular APIs (by decoupling computation and data dependencies, and enabling seamless integration with existing LLM frameworks, such as PyTorch FSDP, Megatron-LM and vLLM), and flexible device mapping (supporting various placement of models onto different sets of GPUs for efficient resource utilization and scalability across different cluster sizes). veRL is fast with state-of-the-art throughput and efficient actor model resharding with 3D-HybridEngine.
DeepScaleR releases DeepScaleR-1.5B-Preview, a 1.5B model, trained on 40K high-quality math problems with 3,800 A100 hours, outperforming OpenAI’s o1-preview on multiple competition-level math benchmarks (such as surpassing O1-Preview and achieves 43.1% Pass@1 on AIME). It is achieved by iteratively scaling Deepseek’s GRPO algorithm from 8K->16K->24K context length for thinking. It releases the training recipe, hyperparameters, data and underlying systems. For the training dataset, it compiled AIME problems and AMC problems prior to 2023, along with questions from the Omni-MATH and Still datasets, which feature problems from various national and international math competitions. The data processing pipeline consists of three key steps: extracting answers, removing redundant questions, and filtering ungradable questions. After deduplication and filtering, the final training dataset consists of approximately 40,000 unique problem-answer pairs.
Different from the original DeepScaleR to train a 1.5B LLM with RL, we adopt the DeepScaleR framework to finetune our 7B DPO model with RL for improving CoT capabilities. As advocated in Deepseek-R1, we employ an Outcome Reward Model (ORM) as opposed to a Process Reward Model (PRM) to avoid reward hacking. The reward function returns: 1 if the LLM’s answer passes basic LaTeX/Sympy checks, and 0 if the LLM’s answer is incorrect or formatted incorrectly. We evaluate our COT reasoning model on competition-level mathematics benchmarks, including AMC 2023, MATH-500, Minerva Math, and OlympiadBench. We report the Pass@1 accuracy. We use Qwen Math’s codebase for evaluation following [109, 110].
Besides DeepScaleR, we adopt another framework, AReal [69, 68] to finetune our DPO model and obtain another version of our COT model. AReaL (Ant Reasoning RL) [69, 68] is an open-source and efficient reinforcement learning system developed at the RL Lab, Ant Research. AReaL inherits and adapts the Open-Source Project ReaLHF [69] for training Large Reasoning Models (LRMs). Its RL training dataset consists of 40k high-quality mathematical reasoning tasks released by DeepScaleR. It assigns rewards of +5 for correct responses and -5 for incorrect responses. The correctness of responses is evaluated by comparing the responses with the answers using Sympy, following the evaluation protocol of Qwen. A simplified variant of PPO is adopted as the RL algorithm with a major change to PPO by eliminating the critic to reduce computation resources. We follow the default configuration of AReal (with necessary modifications due to our specific environment) to train our 7B DPO model for COT reasoning capabilities.